OpenStudy (autogenius):

Physics help. Vector A⃗ points in the negative x direction. Vector B⃗ points at an angle of 27.0∘ above the positive x axis. Vector C⃗ has a magnitude of 10m and points in a direction 40.0∘ below the positive x axis. Given that A⃗ +B⃗ +C⃗ =0, find the magnitudes of A⃗ and B⃗ . Express your answer using two significant figures. Can you please show work?

OpenStudy (danjs):



OpenStudy (danjs):

\[R = \sum F = A + B + C = 0\] \[\sum F _{x} = 0 = A _{x}+B _{x}+C_{x}\] \[\sum F _{y}=0=A _{y}+B _{y}+C _{y}\]

OpenStudy (danjs):

\[A = <-\left| \left| A \right| \right|, 0>\] \[B = <\left| \left| B \right| \right|\cos(27), \left| \left| B \right| \right|\sin(27)>\] \[C = <10\cos(40), -10\sin(40)>\]

OpenStudy (danjs):

Set up two equations, sum of x direction forces, and sum of y direction forces, both equal zero since it is in equilibrium. The two unknown values are the magnitudes of vectors A and B.

OpenStudy (autogenius):

Thank you so much!